An aqueous ideal solution of liquid X (molecular mass56)28%byweight has vapour pressure of 150mm. Find the vapour pressure of X of vapour pressure of water is155mm Hg.

Grade:12th pass

1 Answers

Deepak Sharma
5 years ago
No of moles of X=28/56=0.5
No. of moles of water=72/18=4
Db=0.5/(0.5+4)=1/9
Ca=1-1/9=8/9
Pb=p°b×(xb)=p°b×1/9
Pa=p°a×(xa)=p°a×8/9
P=pa+pb
150=155×8/9+p°b×1/9
So p°b=109.8mm.